5) The displacement, 𝑦(m), of a body in damped oscillation is

𝒚 = 𝟐𝒆 −𝒕 𝐬𝐢𝐧 𝟑𝒕

a) Use the Product Rule to find an equation for the velocity of the object if 𝑣 = 𝑑𝑦 𝑑𝑡


6) You are given the following function which describes the readout taken from an oscilloscope

𝑦 = 3cos (2𝑥)

Your tasks are to:

a) Create a graph showing 2 full periods of the function

b) Take two gradients from the graph

c) Differentiate the function and prove your answers from part b using calculus

d) Compare your answers 


1
Expert's answer
2021-11-07T16:01:47-0500

5)

"v(t)=y'(t)=-2e^{-t}sin3t+6e^{-t}cos3t"


6)

a)



b)

gradients from the graph:

at point (1.1, -2):

"\Delta y/\Delta x=-1.5/0.4=-3.75"

at point (5.1, -2):

"\Delta y/\Delta x=1.4/0.4=3.5"


c)

"y'=-6sin(2x)"

"y'(1.1)=-4.85"

"y'(5.1)=4.20"


d)

error at point (1.1, -2):

"\frac{4.85-3.75}{4.85}=22.7\%"


error at point (5.1, -2):

"\frac{4.2-3.5}{4.2}=16.67\%"


to compare results it was used formula for relative error:


"\frac{|gr_t-gr_g|}{gr_t}"


where grt is gradient from theoretical calculations,

grr is gradient from graph